OpenStudy (ammarah):

Solve the equation by finding the square roots. 1. 4x^2 -4x+1=36 2. x^2 -2/3x+1/9=1 3. 9x^2 +12x +4 =5

OpenStudy (mathstudent55):

1. \(4x^2 -4x+1=36\) First, can you factor the left side? Hint: It's the square of a binomial.

OpenStudy (ammarah):

wait is it x^2 -x +1/4 =9??

OpenStudy (mathstudent55):

No. Just take \(4x^2 - 4x + 1 \) and factor it into the square of a binomial.

OpenStudy (mathstudent55):

What you wrote is correct, but it doesn't help.

OpenStudy (ammarah):

i dont get it what do u mean by square binomia.ll is it 4x^ -2x +1?

OpenStudy (mathstudent55):

Haven't you learned how to factor trinomials? For example, \(x^2 + 3x + 2\) factors into \( (x + 2)(x + 1) \).

OpenStudy (ammarah):

but i dk how to factor this one cuz it has a coefficient infront of the x^

OpenStudy (ammarah):

can u please show me how to do this one....

OpenStudy (mathstudent55):

Ok, here it is. To factor a trinomial of the form ax^2 + bx + c, start by 1. multiplying a and c together. 2. Then find 2 numbers that multiply to ac and add to b. Let's call them m and n. 3. Then break up the middle term of the trinomial into bx = mx + nx. 4. Then factor the polynomial ax + mx + nx + c by grouping. That means, factor a common factor out of the first two terms, and another common factor out of the last 2 terms. 5. Factor out the common factor.

OpenStudy (mathstudent55):

Let's use your first problem as an example.

OpenStudy (ammarah):

ok...

OpenStudy (mathstudent55):

\(4x^2 -4x+1=36\) For now we just want to factor the left side.

OpenStudy (ammarah):

yea m having trouble with that

OpenStudy (ammarah):

wait a second can i dived by 4 to get rid of the coefficient before x^2???

OpenStudy (mathstudent55):

That's why we're going to do it together. Compared to \(\color{red}{a}x^2 + \color{green}{b}x + \color{blue}{c}\), you see that for the trinomial \(\color{red}{4}x^2 \color{green}{- 4}x + \color{blue}{1}\), \(\color{red}{a = 4}\), \(\color{green}{b = -4}\) and \(\color{blue}{c = 1}\). (I hope the colors make it clear.)

OpenStudy (mathstudent55):

No, don't divide by 4.

OpenStudy (mathstudent55):

Since we have a = 4, and c = 1, we multiply ac to get ac = 4(1) = 4.

OpenStudy (mathstudent55):

Now we need two factors of 4 that add to b, -4.

OpenStudy (mathstudent55):

Those would be -2 and -2, since (-2)(-2) = 4, and (-2) + (-2) = -4.

OpenStudy (mathstudent55):

Now we break up the middle term into those two numbers we found: \(4x^2 - 4x + 1\) \( = 4x^2 - 2x - 2x + 1 \) Now we factor by grouping.

OpenStudy (ammarah):

so the answer would be x =8 or 4

OpenStudy (mathstudent55):

I have no idea about the answer. We are still factoring. \(4x^2 - 2x - 2x + 1 \) \(=2x(2x - 1) - 1(2x - 1) \) Now we see that the line above has a common factor. See below in color: \(=\color{blue}{2x}\color{red}{(2x - 1)} \color{green}{- 1}\color{red}{(2x - 1)} \) Now we factor out the red common factor to find the factorization: \( \color{red}{(2x - 1)}(\color{blue}{2x} \color{green}{- 1}) \) The final factorization is: \( (2x - 1)^2 \)

OpenStudy (mathstudent55):

Now that the left side is factored, that means we have \( (2x - 1)^2 = 36\) In a previous post of yours, I wrote that when you have an equation of the form \(x^2 = b\), you solve it by doing the following: \(X = \pm \sqrt{b} \) Now we apply this here: \( 2x - 1 = \pm \sqrt{36} \) \(2x - 1 = 6\) or \(2x - 1 = -6 \) \(2x = 7\) or \(2x = -5\) \(x = \dfrac{7}{2} \) or \(x = -\dfrac{5}{2} \)

OpenStudy (ammarah):

ok i get it thanks for the next one how would u factor it?

OpenStudy (mathstudent55):

\(x^2 -\dfrac{2}{3}x+\dfrac{1}{9}=1 \) My suggestion for the second one is to first multiply both sides by 9 to get rid of all denominators.

OpenStudy (ammarah):

@ganeshie8

OpenStudy (mathstudent55):

\(9(x^2 -\dfrac{2}{3}x+\dfrac{1}{9})=9(1)\) \(9x^2 -6x+1=9\) \((3x - 1)^2 = 9 \) \( 3x - 1 = \pm\sqrt{9} \) \(3x - 1 = 3\) or \(3x - 1 = -3\) 3x = 4 or 3x = -2 \(x = \dfrac{4}{3}\) or \(x = -\dfrac{2}{3} \)